# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulation for Time-Efficient Fine-Resolution Policy Learning

### v1

@@ I Introduction
Autonomous excavation has been studied to meet the rising demand for earthwork and construction [1]. In excavation, especially in mountainous areas, quarry sites, and construction sites, large rocks are often mixed with sand, gravel, and pebbles in the soil [2]. Operating excavators must have advanced operational expertise to find, collect, and transport these rocks [3]. For rock excavation, excavators must efficiently lift and move rocks using a bucket. This involves non-grasping bucket motion while considering forces from rock-soil-bucket interactions, a task only skilled operators can perform [4]. Due to the varied shapes and sizes of rocks, automating rock excavation with simple autonomous control policies is challenging [5].
@@
Reinforcement learning (RL) is potentially useful to learn optimal actions from diverse state-action transitions caused by various rocks [6]. RL learns control policies from interaction samples between a robot and the environment. Thus, simulators are commonly used since real-world sampling requires significant manual efforts to set up various types of soil and objects [7]. In previous simulation work, computational mechanics and particle methods have been used. Computational mechanics, which discretize terrain into elements, are suitable for simple terrains but struggle with complex terrains involving rocks or nonlinear phenomena [7]. Conversely, particle methods model materials as individual particles and simulate their interactions, effectively reproducing complex phenomena like friction, collisions, and soil deformation [8]. This makes particle methods suitable for simulating complex terrains that include rocks.
While particle-based simulations can be accurate and realistic, they face challenges in calculation time for our RL application, particularly as spatial resolution improves. This resolution can be defined by the particle size/number in the whole soil space, with smaller particle sizes (more amounts of particles) allowing for a more detailed particle representation of the soil space and obtaining a finer resolution. Fine resolution leads to complex particle interactions, making it impractical to collect the vast samples needed for RL in a reasonable calculation time [9]. Ideally, the resolution should be coarsened until the simulator is sufficiently fast; however, simulation deviates from real-world behavior, utilizing the learned policies ineffective in real-world environments.
@@
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) for achieving time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. PRPD progressively improves simulation resolution while learning and transferring policies at each stage. PRPD uses a conservative policy transfer scheme to regularize policies, stabilizing policy transfer across different resolution simulations having different behaviors. We demonstrated the effectiveness of PRPD by constructing a variable-resolution rock excavation simulator. PRPD reduced sampling time to less than 1/7 while maintaining task success rates comparable to policy learning in a fine-resolution simulation. Furthermore, in evaluations of the learned policies within a real-world environment containing nine types of rocks, PRPD achieved rock excavation in approximately 90 % of scenarios. As an impact of this study, achieving realistic simulations of complex environments, such as excavation tasks, remains challenging. While advances in computational resources and simulation technologies may alleviate simulation time issues, simulating real-world environments accurately remains prohibitively costly and will still face resource limits. This study offers insights into addressing these challenges by improving policy-learning efficiency.
@@ II-A Learning Excavation Policy
This section describes several methods for acquiring control policies for robotic excavations with soil and rocks. Previous works have proposed the following three approaches.
Learning in Real-world: Learning rock-excavation tasks in real-world environments requires extensive sampling time. These works have modeled soil and rock behavior from limited samples to learn control policies [5]. However, accurately modeling rock behavior requires interaction samples for various rock shapes. The vast number of combinations makes it difficult to apply this method to multiple shapes due to sampling costs, so previous research has only obtained policies for a single rock shape.
Learning in Simulation by Computational Mechanics: These works simulate soil behavior in response to bucket movements using computational mechanics and collect learning samples from model interactions [7]. The finite element method divides soil into small elements and solves for displacement, stress, and strain. Deformation analysis uses soil mechanics equations to analyze deformation and failure. These works are suitable for simulating simple terrain because they numerically discretize and simulate terrain, but they have difficulty accurately reproducing complex terrain involving rocks and nonlinear phenomena.
Learning in Simulation by Particle Method: These works adopt the particle method that models materials as individual particles and directly simulates the interactions between them [10]. This enables detailed simulation of rock and soil behavior, and approximates complex physical phenomena such as inter-particle friction, collision, and soil deformation separately. However, calculating interactions between multiple particles is time-consuming, making it difficult to collect many learning samples.
We aim to establish learning policies applicable to various rock shapes by utilizing a particle-based simulator capable of achieving this aim. To address the sampling time challenge, we propose a novel policy learning framework that enables the short-time learning sample collection from various rock shapes, thereby achieving multiple-shape rock excavation.
@@ II-B Excavation Simulator by Particle Method
Excavation simulators with a large number of particle interactions require extensive calculation time [10]. The following paragraphs describe previous works that employed three approaches to accelerate calculations and comparisons with our developed simulator.
Static Adjustment of Resolution: To reduce the calculation time for particle interactions, which make up the majority of simulation calculations, efforts have focused on generating environments limited to specific work areas and approximating many soil particles with fewer macro particles [11]. These strategies have decreased the necessary memory resources and sampling time. However, there are limits to how much the resolution can be reduced without diverging from real-world behavior.
Dynamic Adjustment of Resolution: Some works accelerate simulations by dynamically changing the resolution [8]. They dynamically estimate the work area, splitting particles in fine-resolution regions and merging particles in coarse-resolution areas, thereby reducing calculation time. While this approach achieves acceleration, focusing fine resolution only in the work area still requires significant resources and calculation time.
Parallel Calculation of Particles: Recently, frameworks like Isaac Gym and Isaac Sim have been developed to rapidly simulate robot environments using GPUs, significantly accelerating sampling time [11]. In excavation environments, these achieve faster calculations by parallelizing the calculation of multiple particle interactions. However, generating a vast number of particles within the simulation requires substantial GPU memory, making it difficult to create parallel environments that can further reduce sampling time [12].
Our Developed Simulator: The simulator developed in this paper incorporates equivalents of the three acceleration techniques of the previous works: (1) approximating soil with micro particles, (2) generating only the work area for fine-resolution simulation, and (3) using Isaac Gym for GPU acceleration. The details of our simulator are described in Sec. V-D.
Despite these advancements in simulation acceleration, they remain insufficient for the vast sample collection required in RL [8]. The goal of this paper is not to develop realistic simulations but to learn policies of realistic simulations. To shorten sampling time, we use both fine-resolution simulations that take longer to calculate and extremely coarse-resolution simulations that require less calculation time.
@@ II-C Sample Efficient Reinforcement Learning Methods
Several methods improve sample efficiency in RL. Model-based RL uses a learned model of the environment’s dynamics [13]. Auxiliary tasks and autoencoders extract features from observations [13, 14]. Data augmentation enhances information content by transforming observation samples [15]. Ensemble learning improves robustness and adaptability with multiple models [14]. Imitation learning trains policies from demonstrations, and reward function learning from small demonstration sets reduces human effort [9]. Off-policy RL such as soft actor-critic (SAC) efficiently explores and learns by reusing past interaction data [16].
Various methods have improved RL’s “sample efficiency” by extracting more information from each sample or increasing update times to maximize sample utilization. However, these approaches often come with computational overhead, such as increased calculation time and resource demands by adding learning components, which can compromise “time efficiency,” the overall training time for policy learning.
Our approach can be interpreted as a type of curriculum reinforcement learning (CRL) since it also involves gradually changing the tasks from simple to complex [17]. In detail, CRL gradually increases task difficulty, reducing the sample requirements compared to learning solely on complex tasks. However, to the best of our knowledge, our approach is the first to treat differences in the spatial resolution of complex simulations as distinct tasks.
@@ III-A Reinforcement Learning
Reinforcement Learning (RL) is formulated as a Markov Decision Process (MDP) $\mathcal{M}$. MDP is defined by the state space $\mathcal{S}$, action space $\mathcal{A}$, state transition probability $P(s^{\prime}|s,a)$, and reward function $R(s,a)$, represented as $\mathcal{M}\!=\!(\mathcal{S},\mathcal{A},P,R)$. The agent selects an action $a\in\mathcal{A}$ in the current state $s\in\mathcal{S}$ and, as a result, observes the next state $s^{\prime}\in\mathcal{S}$ and reward $r\in\mathbb{R}$. The goal in an MDP is to find a policy $\pi(a|s)$ that maximizes the expected cumulative reward $J(\pi)$, given by $J(\pi)\!=\!\mathbb{E}_{\pi}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}r_{t}\right]$, $\gamma\!\in\![0,1]$ is the discount factor.
To maximize the expected cumulative reward $J(\pi)$, policy gradient methods can be used, where the objective function is defined as $J(\pi)=\mathbb{E}_{\pi}\left[\sum\nolimits_{t=0}^{\infty}A(s_{t},a_{t})\right]$, where $A(s,a)=Q(s,a)-V(s)$ is the advantage function. The action value function $Q(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s,a_{0}=a\right]$ represents the expected cumulative reward when action $a$ is taken in state $s$. The state value function $V(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s\right]$ represents the expected cumulative reward starting from state $s$ and following policy $\pi(s,a)$. These functions are typically approximated using function approximators, specifically neural networks, which parameterize the policy as $\pi(s,a;\theta)$, the state value function as $V(s;\phi)$, and the action value function as $Q(s,a;\psi)$, where $\theta$, $\phi$, and $\psi$ represent their respective parameters.
Typically, RL methods learn an optimal policy by iteratively conducting the following two steps: (1) collecting samples $(s,a,r,s^{\prime})$ with the policy $\pi$ into a rollout buffer $\mathcal{D}_{i}$, and (2) updating the policy using these samples. This iteration number is denoted as $i$, and the parameter update number is denoted as $k$.
@@ III-B Conservative Policy Update in Reinforcement Learning
In RL, there are various errors such as function approximation errors and observation noise, and attempts have been made to learn policies that are robust to these errors. Conservative Policy Iteration (CPI) is one example, which is designed to achieve more stable policy updates by relaxing the policy updates scheme [18].

### v2

# Progressive-Resolution Policy Distillation: Leveraging Coarse-Resolution Simulations for Time-Efficient Fine-Resolution Policy Learning

@@ Abstract
Note to Practitioners
This paper is motivated by the issue of computation time in excavation simulation using soil particles. The behavior of real soil is highly complex, and approximating it at high resolution requires enormous computational costs. Therefore, existing soil simulators have focused on improving simulation accuracy while maintaining reduced computation time. This paper takes a different approach by focusing on the learning of control policies in excavation simulators and proposes a framework for reducing calculation time in such use cases. In this framework, a control policy is first learned in a low-resolution simulation, significantly reducing computation time. The learned policy is then transferred to a high-resolution simulation for retraining, thereby achieving an overall reduction in simulation time. Furthermore, to enable robust policy transfer across different resolutions, this paper discusses a stable policy distillation scheme and insights into resolution design. This approach enables the development of autonomous excavation systems without relying on expensive real-world data collection, improving the scalability and adaptability of autonomous excavation. Simulation experiments suggest that this framework significantly reduces training time compared to conventional policy learning approaches. However, real-world validation has so far been limited to simple excavation robots. Future research will explore applications to excavators and other machinery more suitable for real-world operations. Although this paper focuses on autonomous excavation, the proposed approach can also be extended to environments where increased simulation resolution critically impacts computation time, such as liquid and soft object manipulation.
Index Terms:
Simulation Resolution, Excavation, Policy Distillation, Reinforcement Learning
@@ I Introduction
Autonomous excavation has been studied to meet the rising demand for earthwork and construction [1, 2, 3]. In excavation, especially in mountainous areas, quarry sites, and construction sites, large rocks are often mixed with sand, gravel, and pebbles in the soil [4, 5]. Operating excavators must have advanced operational expertise to find, collect, and transport these rocks [6, 7]. For rock excavation, excavators must efficiently lift and move rocks using a bucket. This involves non-grasping bucket motion while considering forces from rock-soil-bucket interactions, a task only skilled operators can perform [8, 9]. Due to the varied shapes and sizes of rocks, automating rock excavation with simple autonomous control policies is challenging [10, 11].
@@
Reinforcement learning (RL) is potentially useful to learn optimal actions from diverse state-action transitions caused by various rocks [12, 13, 14]. RL learns control policies from interaction samples between a robot and the environment. Thus, simulators are commonly used since real-world sampling requires significant manual efforts to set up various types of soil and objects [15, 16]. In previous simulation work, computational mechanics and particle methods have been used. Computational mechanics, which discretize terrain into elements, are suitable for simple terrains but struggle with complex terrains involving rocks or nonlinear phenomena [15, 16]. Conversely, particle methods model materials as individual particles and simulate their interactions, effectively reproducing complex phenomena like friction, collisions, and soil deformation [17, 18]. This makes particle methods suitable for simulating complex terrains that include rocks.
While particle-based simulations can be accurate and realistic, they face challenges in computation time for our RL application, particularly as spatial resolution improves. This resolution can be defined by the particle size/number in the whole soil space, with smaller particle sizes (more amounts of particles) allowing for a more detailed particle representation of the soil space and obtaining a finer resolution. Fine resolution leads to complex particle interactions, making it impractical to collect the vast samples needed for RL in a reasonable calculation time [19, 20]. Ideally, the resolution should be coarsened until the simulator is sufficiently fast; however, simulation deviates from real-world behavior, utilizing the learned policies ineffective in real-world environments. Our focus is on addressing RL problems where there is a trade-off between resolution-dependent calculation time and sample quality, aiming to reduce the computation time required for policy learning.
@@
We propose a novel policy learning framework called Progressive-Resolution Policy Distillation (PRPD) for achieving time-efficient policy learning of fine-resolution simulations by utilizing coarse-resolution simulations effectively. PRPD progressively improves simulation resolution while learning and transferring policies at each stage. PRPD uses a conservative policy transfer scheme to regularize policies, stabilizing policy transfer across different resolution simulations having different behaviors. We demonstrated the effectiveness of PRPD by constructing a variable-resolution rock excavation simulator using Isaac Gym [21]. This paper demonstrated a 7-fold improvement in learning time efficiency (PRPD: 90 minutes, policy learning in the finest-resolution simulation: 600 minutes), resulting in an 8-hour time gap. This difference is likely to become more significant with increasing state-action space complexity or task difficulty. Furthermore, in evaluations of the learned policies within a real-world environment containing nine types of rocks, PRPD achieved rock excavation in approximately 90% of scenarios.
As an impact of this study, achieving realistic simulations of complex environments, such as excavation tasks, remains challenging. While advances in computational resources and simulation technologies may alleviate simulation time issues, simulating real-world environments accurately remains prohibitively costly and will still face resource limits. This study offers insights into addressing these challenges by improving policy-learning efficiency.
@@ II-A Learning Excavation Policy
This section describes several methods for acquiring control policies for a robotic excavation task with soil and rocks. Previous works have proposed the following three approaches.
Learning in Real-world: Learning rock-excavation tasks in real-world environments requires extensive sampling time. These works have modeled soil and rock behavior from limited samples to learn control policies [10, 1]. However, accurately modeling rock behavior requires interaction samples for various rock shapes. The vast number of combinations makes it difficult to apply this method to multiple shapes due to sampling costs, so previous research has only obtained policies for a single rock shape.
Learning in Simulation by Computational Mechanics: These works simulate soil behavior in response to bucket movements using computational mechanics and collect learning samples from model interactions [15, 16]. The finite element method divides soil into small elements and solves for displacement, stress, and strain. Deformation analysis uses soil mechanics equations to analyze deformation and failure. These works are suitable for simulating simple terrain because they numerically discretize and simulate terrain, but they have difficulty accurately reproducing complex terrain involving rocks and nonlinear phenomena.
Learning in Simulation by Particle Method: These works adopt the particle method that models materials as individual particles and directly simulates the interactions between them [22, 23]. This enables detailed simulation of rock and soil behavior, and approximates complex physical phenomena such as inter-particle friction, collision, and soil deformation separately. However, calculating interactions between multiple particles is time-consuming, making it difficult to collect many learning samples. While imitation learning applications exist with few samples, these methods have not been trained in comprehensive environments and can only be applied in limited situations.
This paper aims to establish learning policies applicable to various rock shapes by utilizing a particle-based simulator capable of achieving this aim. To address the sampling time challenge, we propose a novel policy learning framework that enables the short-time collection of learning samples from various rock shapes, thereby achieving rock excavation for multiple shapes.
@@ II-B Excavation Simulator by Particle Method
Excavation simulators with a large number of particle interactions require extensive calculation time [22, 18]. The following paragraphs describe previous works that employed three approaches to accelerate calculations and comparisons with our developed simulator.
Static Adjustment of Resolution: To reduce the calculation time for particle interactions, which make up the majority of simulation calculations, efforts have focused on generating environments limited to specific work areas and approximating many soil particles with fewer macro particles [20, 24]. These strategies have decreased the necessary memory resources and sampling time. However, there are limits to how much the resolution can be reduced without diverging from real-world behavior.
Dynamic Adjustment of Resolution: Some works accelerate simulations by dynamically changing the resolution [17, 19]. They dynamically estimate the work area, splitting particles in fine-resolution regions and merging particles in coarse-resolution areas, thereby reducing calculation time. While this approach achieves acceleration, focusing fine resolution only in the work area still requires significant resources and calculation time, making it unsuitable for the extensive sample collection needed in RL.
Parallel Calculation of Particles: Recently, frameworks like Isaac Gym and Isaac Sim have been developed to rapidly simulate robot environments using GPUs, significantly accelerating sampling time [21, 25]. In excavation environments, these achieve faster calculations by parallelizing the calculation of multiple particle interactions. However, generating a vast number of particles within the simulation requires substantial GPU memory, making it difficult to create parallel environments that can further reduce sampling time [24].
Our Developed Simulator: The simulator developed in this paper incorporates equivalents of the three acceleration techniques of the previous works: (1) approximating soil with micro particles, (2) generating only the work area for fine-resolution simulation, and (3) using Isaac Gym for GPU acceleration. The details of our simulator are described in Section V-D.
Despite these advancements in simulation acceleration, they remain insufficient for the vast sample collection required in RL [17, 19]. RL needs millions of samples, making traditional simulators challenging for policy learning [12, 13]. The goal of this paper is not to develop realistic simulations but to learn policies of realistic simulations. To shorten sampling time, we use both fine-resolution simulations that take longer to calculate and extremely coarse-resolution simulations that require less calculation time.
@@ II-C Sample Efficient Reinforcement Learning Methods
Several methods improve sample efficiency in RL. Model-based RL uses a learned model of the environment’s dynamics [26, 27, 28]. Auxiliary tasks, autoencoders, and contrastive learning extract better features from observations [29, 30, 31]. Data augmentation enhances information content by transforming observation samples [32, 33, 34]. Ensemble learning improves robustness and adaptability with multiple models [31, 35, 36]. Imitation learning trains policies from demonstrations, and reward function learning from small demonstration sets reduces human effort [37, 38, 39]. Off-policy RL such as soft actor-critic (SAC) efficiently explores and learns by reusing past interaction data [40, 41].
Various methods have improved RL’s “sample efficiency” by extracting more information from each sample or increasing update times to maximize sample utilization. However, these approaches often come with computational overhead, such as increased calculation time and resource demands by adding learning components, which can compromise “time efficiency,” the overall training time for policy learning. Since this paper focuses on time efficiency, our experiments utilize on-policy RL, which provides high time efficiency due to its simpler learning structure, fewer model parameters, and reduced update frequency [42, 43].
Our approach can be interpreted as a type of curriculum reinforcement learning (CRL) since it also involves gradually changing the tasks from simple to complex [44, 45]. In detail, CRL gradually increases task difficulty, reducing the sample requirements compared to learning solely on complex tasks. However, to the best of our knowledge, our approach is the first to treat differences in the spatial resolution of complex simulations as distinct tasks.
@@ III-A Reinforcement Learning
Reinforcement Learning (RL) is a framework where an agent learns an optimal policy through interaction with the environment [46]. Typically, RL is formulated as a Markov Decision Process (MDP) $\mathcal{M}$. An MDP is defined by the state space $\mathcal{S}$, action space $\mathcal{A}$, state transition probability $P(s^{\prime}|s,a)$, and reward function $R(s,a)$, represented as $\mathcal{M}=(\mathcal{S},\mathcal{A},P,R)$. The agent selects an action $a\in\mathcal{A}$ in the current state $s\in\mathcal{S}$ and, as a result, observes the next state $s^{\prime}\in\mathcal{S}$ and reward $r\in\mathbb{R}$. The goal in an MDP is to find a policy $\pi(a|s)$ that maximizes the expected cumulative reward $J(\pi)$, given by:
$J(\pi)=\mathbb{E}_{\pi}\left[\sum\nolimits_{t=0}^{\infty}\gamma^{t}r_{t}\right],$ (1)
where $\gamma\in[0,1]$ is the discount factor.
To maximize the expected cumulative reward $J(\pi)$, policy gradient methods are used, where the objective function is defined as:
$J(\pi)=\mathbb{E}_{\pi}\left[\sum\nolimits_{t=0}^{\infty}A(s_{t},a_{t})\right],$ (2)
where $A(s,a)=Q(s,a)-V(s)$ is the advantage function. The value function $Q(s,a;\psi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s,a_{0}=a\right]$ represents the expected cumulative reward when action $a$ is taken in state $s$. The value function $V(s;\phi)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{t}\big{|}s_{0}=s\right]$ represents the expected cumulative reward starting from state $s$ and following policy $\pi(s,a;\theta)$. These functions are typically approximated using function approximators such as neural networks. Here, $\theta$, $\phi$, and $\psi$ denote the parameters of the policy $\pi$, the value functions $V,Q$, respectively.
A representative method based on policy gradient methods is proximal policy optimization (PPO) [42]. PPO learns an optimal policy by iteratively collecting samples $(s,a,r,s^{\prime})$ with the policy $\pi$ into a rollout buffer $\mathcal{D}$, and updating the policy using these samples. The number of such iterations is denoted as $i$, and the number of parameter updates is denoted as $k$. PPO updates the policy by minimizing the following loss function:
$\begin{cases}L_{\theta,\phi}^{\pi,V}\!=\!\!\!&\mathbb{E}_{\mathcal{D}_{i}}[\min\{\rho A_{k}(s,a),\text{clip}(\rho,1-\epsilon,1+\epsilon)A_{k}(s,a)\}\\ &-c_{1}(V_{k}(s)-\hat{R})^{2}+c_{2}\pi_{k}(a|s)\log\pi_{k}(a|s)],\end{cases}$ (3)
where $\rho$ represents the importance sampling ratio, and $\epsilon$ is a hyperparameter controlling the clipping range. The parameters $c_{1}$ and $c_{2}$ control the weights of the value function loss and the entropy regularization, respectively. $\hat{R}$ is the cumulative reward calculated from the samples in $\mathcal{D}$ [42].
@@ III-B Conservative Policy Update in Reinforcement Learning
In RL, there are various errors such as function approximation errors and observation noise, and attempts have been made to learn policies that are robust to these errors. Conservative Policy Iteration (CPI) is one example, which is designed to achieve more stable policy updates by relaxing the policy updates scheme [47, 48].
@@
Algorithm 1 Progressive-Resolution Policy Distillation with PPO # Set parameters described in Table I
# Schedule resolution progressively $(\Delta_{1},\Delta_{2},...,\Delta_{N})$
for $n=1,2,...,N$ do
       # Initialize iteration num $i\!=\!0$, parameter update num $k\!=\!0$
       # Set values $V^{(n)}_{k}$, $Q^{(n)}_{k}$, policy $\pi^{(n)}_{k}$, rollout buffer $\mathcal{D}^{(n)}_{i}$
       # Copy network parameters ($\Delta_{n-1}$ to $\Delta_{n}$ )
       # Set resolution $\Delta\leftarrow\Delta_{n}$ and generate simulator $\mathcal{M}^{(n)}$
       while until $\tau>\hat{\tau}$ do
             for $e=1,2,...,E$ do
                   for $t=1,2,...,T$ do
                         # Take action $a_{t}$ from $\pi^{(n)}_{k}$ in $\mathcal{M}^{(n)}$
                         # Get observation $s_{t}$, reward $r_{t}$
                         # Push $(s_{t},a_{t},r_{t})$ to $\mathcal{D}^{(n)}_{i}$
            for $k^{\prime}=k,k+1,...,k+K$ do
                   # Calculate loss: Eq. (10), update $V^{(n)}_{k},Q^{(n)}_{k},\pi^{(n)}_{k}$
            # update numbers $i=i+1$, $k=k+K$
             # Check success rate $\tau$
TABLE I: Learning parameters of PRPD in experiments.
Para. Meaning Value
$\alpha_{0}$ Scaling coefficient of distillation 2
$\gamma$ Discount factor of RL 0.99
$\Delta_{1}$ Scale of initial resolution [mm] 70
$\Delta_{N}$ Scale of final resolution [mm] 10
$\Delta_{\mathcal{R}}$ Scale of resolution interval [mm] 10
$\hat{\tau}$ Target success rate 0.95
$T$ Number of steps per episode 128
$E$ Number of episodes per iteration 128
$K$ Number of parameter update per iteration 64$\times$8